Circular No. 8005

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002hh IN NGC 6946 Further to IAUC 7906, W. Li reports the LOTOSS discovery of an

apparent supernova on unfiltered KAIT images taken on Oct. 31.1 (mag ∼ 16.5) and Nov. 2.1 UT (mag ∼ 16.3). The new object is located at α = 20^{h}34^{m}44^{s}.29, δ = +60^{o}07′19′′.0 (equinox 2000.0), which is 60′′.9 west and 114′′.1 south of the nucleus of NGC 6946. A KAIT image taken on Oct. 26.1 showed nothing at this position (limiting mag ∼ 19.0). NGC 6946 has produced six known supernovae: 1917A, 1939C, 1948B, 1968D, 1969P, and 1980K. Observations of this potentially bright supernova are encouraged.

V838 MONOCEROTIS U. Munari and S. Desidera, Asiago Astrophysical Observatory; and A.

Henden, Universities Space Research Association and U.S. Naval Observatory (USNO), communicate: “Photometry obtained with the USNO 1.0-m telescope shows the central star of V838 Mon to be constant in the blue around V = 16.05, U − B = −0.07, and B − V = +0.69 since September, while brightening by 0.012 and 0.020 mag/day in bands R and I, respectively (from R = 14.43 and I = 10.93 on Sept. 25.5 UT to R = 14.00 and I = 10.22 on Oct. 30.5). Spectroscopic observations (range 360–774 nm at 0.32 nm/pixel) of V838 Mon were obtained on Oct. 28 and 29 with the Asiago 1.82-m telescope (+ AFOSC). The hot companion star reported on IAUC 7982 (confirmation report on IAUC 7992) seems to resemble a normal B3 V star, suffering from a reddening of E(B − V) = +0.9 that is consistent with a law having R_{V} = A_{V}/E(B − V) = 3.1 (cf. Savage and Mathis 1979, ARAA 17, 73), in agreement with earlier estimates based on the intensity of Na I and K I interstellar lines (cf. IAUC 7812). However, if the companion is a genuine B3 V star, V838 Mon lies 10500 pc distant and just 200 pc above the galactic plane, for which such a low reddening is surprising, and the mass of the M-giant progenitor should have been larger than 7 M_{⊙}. The fact that the B3 V star is ∼ 1 mag fainter in B than V838 Mon in quiescence, and that the optical-infrared energy distribution of the latter closely matched an F0 star (luminosity class III-II for a distance of 10500 pc), suggests that the quiescent light could have been dominated by the progenitor of the outbursting companion or by circumstellar material in the form of a warm disk, or something else. The current spectrum of the cool component holds identical to the ‘later-than-M10’ classification of one month ago (cf. IAUC 7982).”

2002 November 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green